OpenStudy (anonymous):

What is the solution of the equation? sqrt(2x+13)-5=x

OpenStudy (anonymous):

\[\sqrt{2x+13}-5=x\]

OpenStudy (solomonzelman):

\[\sqrt{2x+13}-5=x\]\[\sqrt{2x+13}=x+5~~~~~~~~squ are~~~both~~sides,\]\[2x+13=x^2+10x+25~~~~~~~~Can~~you~~do~~ i t~~now?\]

OpenStudy (anonymous):

Yes,thank you :)

OpenStudy (solomonzelman):

Anytime!

OpenStudy (anonymous):

I thought i could finish it,but im confused so can you help me?

OpenStudy (solomonzelman):

\[x^2+10x+25-2x^2-13=0\]

OpenStudy (solomonzelman):

\[-x^2+10x+12=0~~~~~~->~~~~~~~~x^2-10x-12=0\]

OpenStudy (solomonzelman):

Which methods do you know? Quadratic formula Completing the square Factoring

OpenStudy (solomonzelman):

All of them?

OpenStudy (anonymous):

i know all of them but mainly quadratic formula

OpenStudy (solomonzelman):

Sorry, the connection snapped

OpenStudy (solomonzelman):

\[\huge\color{blue}{\frac{-(-10)±\sqrt{(-10)^2-4 \times 1 \times (-12)}}{2(1)}=x}\]

OpenStudy (solomonzelman):

\[\huge\color{blue}{\frac{+10±\sqrt{100-(-48)}}{2}=x}\] \[\huge\color{blue}{\frac{+10±\sqrt{148)}}{2}=x}\] \[\huge\color{blue}{\frac{+10±4\sqrt{37)}}{2}=x}\] \[\huge\color{blue}{+5±2\sqrt{37}=x}\]

OpenStudy (solomonzelman):

Clear?

OpenStudy (anonymous):

yes very lol thank you so much